# HyperEvent: A Strong Baseline for Dynamic Link Prediction via Relative Structural Encoding

### v2

@@ Abstract
Learning representations for continuous-time dynamic graphs is critical for dynamic link prediction. While recent methods have become increasingly complex, the field lacks a strong and informative baseline to reliably gauge progress. This paper proposes HyperEvent, a simple approach that captures relative structural patterns in event sequences through an intuitive encoding mechanism. As a straightforward baseline, HyperEvent leverages relative structural encoding to identify meaningful event sequences without complex parameterization. By combining these interpretable features with a lightweight transformer classifier, HyperEvent reframes link prediction as event structure recognition. Despite its simplicity, HyperEvent achieves competitive results across multiple benchmarks, often matching the performance of more complex models. This work demonstrates that effective modeling can be achieved through simple structural encoding, providing a clear reference point for evaluating future advancements. The code for HyperEvent is accessible on GitHub [^1] https://github.com/jianjianGJ/HyperEvent.
I Introduction
@@
Temporal graphs model evolving systems through sequences of interactions (i.e., temporal events), serving as essential abstractions for real-world dynamic phenomena like social networks and transaction systems[1, 2, 3]. These graphs are primarily modeled as discrete-time graph snapshots or continuous-time dynamic graphs (CTDGs)[4]. This work focuses specifically on CTDGs because real-world interactions intrinsically manifest as continuous streams of events, where preserving their exact temporal sequence is critical for capturing fine-grained event relationships. Accurately predicting future events within CTDGs, known as dynamic link prediction, remains a core challenge with significant implications for applications such as real-time recommendation[5, 6] and anomaly detection[7, 8].
@@
Recent years have witnessed notable advances in representation learning for CTDGs. Proposed methods have aimed to address challenges such as capturing complex structural-temporal dependencies, modeling long-range interactions, and improving generalization to unseen nodes. These approaches can be broadly categorized into two paradigms: node-centric and event-centric methods. Node-centric methods maintain and update memory states for each node, using recurrent or attention-based mechanisms to evolve node embeddings over time[9, 10, 11, 12]. Some incorporate neural ODEs to model continuous dynamics[13], or employ neighbor caching to alleviate structural bottlenecks[14]. While often scalable and inductive, these techniques reduce each event to an update of its endpoint embeddings, potentially overlooking higher-order relationships between events. In contrast, event-centric methods treat interaction events as primary entities. Techniques such as CAWN[15] leverage causal anonymous walks to extract temporal motifs, while DyGFormer[16] uses attention mechanisms to implicitly model event group patterns. Several recent methods also highlight that relatively simple architectures—such as MLPs or patched transformers—can achieve strong performance[17, 16].
@@
Among these, EdgeBank[18] stands out as a simple, non-parametric baseline that relies on memorizing and replaying historical interactions. However, its performance has been significantly surpassed by newer learning-based models, diminishing its utility as a meaningful benchmark. This creates a critical gap in the literature: the absence of a strong, simple baseline that can serve as a fundamental reference point for evaluating model efficacy. Without such a baseline, it becomes difficult to distinguish genuine advancements from unnecessary complexity, potentially leading to illusory progress.
@@ II Related Work
Temporal Graph Learning for Dynamic Link Prediction. Dynamic link prediction is an essential task for forecasting future interactions in continuously evolving networks, with critical applications across social networks [19, 20], communication systems [21, 22], traffic prediction [23, 24], and knowledge graph completion [25, 26]. Unlike traditional static graph methods that rely on structural heuristics or node similarity metrics [27, 28, 29, 30, 31], dynamic settings require modeling temporal event streams, capturing complex dependencies among interactions, and generalizing to unseen nodes [32, 33]. Early approaches discretized event streams into snapshot sequences [34, 35, 36, 37] and applied graph neural networks [38, 39] combined with sequence models such as RNNs [40, 41] or Transformers [42]. However, such discretization often leads to information loss and limited expressiveness, motivating a shift toward continuous-time modeling.
@@
Continuous-Time Dynamic Graph Learning. CTDG methods can be broadly categorized into node-centric and event-centric paradigms. Node-centric approaches maintain and update memory states for each node upon every interaction. Seminal works include JODIE [10], which uses coupled RNNs to update user and item representations, and DyRep [11], which incorporates multi-hop propagation. The Temporal Graph Network (TGN) framework unifies various memory modules and graph aggregation operators[9], and TGAT employs self-attention to aggregate temporal neighbor information[43]. Though later methods such as TNCN [12] and EDGE [44] improve scalability through asynchronous updates or graph partitioning, they remain constrained by sequential, event-by-event processing that limits parallelism and overlooks higher-order event semantics. On the other hand, event-centric methods explicitly model interactions between events. CAWN [15] and TPNet [45] encode anonymous walks to capture temporal motifs. NAT introduces reusable neighbor representations[14], and DyGFormer applies Transformers to patched event sequences[16], implicitly modeling event group patterns through attention mechanisms. TCL constructs temporal dependency graphs but relies on RNNs that struggle with long-range dependencies[46]. Neural ODE-based approaches like CTAN enable continuous dynamics modeling but often ignore fine-grained event correlations[13]. A common trend across these lines of research is the increasing architectural complexity, with many recent models incorporating sophisticated mechanisms such as attention, memory, or neural differential equations. Despite reported improvements, the field lacks a simple, yet effective baseline that clarifies the minimal complexity required for competitive performance, making it difficult to assess whether proposed enhancements genuinely capture essential temporal-topological phenomena or merely introduce inductive biases.
@@
The Role of Baselines and the Limitations of EdgeBank. Strong baselines are fundamental to meaningful scientific progress in machine learning. They serve as essential references that help the community discern actual advancements from false positives, ensure fair comparisons, and maintain a clear lower bound of performance that new methods should exceed. A poorly performing baseline can misleadingly inflate the perceived utility of new models and encourage unnecessary complexity. In temporal graph learning, EdgeBank has been adopted as a canonical baseline—a simple non-parametric method that predicts future edges based on their recurrence in historical event sequences[18]. While computationally efficient and easy to implement, its performance has fallen significantly behind modern learned models across multiple datasets. This growing performance gap diminishes its utility as a meaningful reference, raising concerns about pseudo-progress in the field. When a simple heuristic is substantially outperformed, it becomes unclear whether newly proposed architectures are capturing genuine temporal dynamics or simply leveraging additional parameters and complexity to memorize dataset-specific patterns. Therefore, there is a pressing need for a new, stronger, and principled baseline that is both conceptually simple and highly competitive, enabling more rigorous evaluation and providing clearer insights into what constitutes necessary model complexity for dynamic link prediction.
III Method
III-A Preliminaries and Problem Formulation
@@
III-B Existing Frameworks
III-B1 Node-centric methods
@@
III-B2 Event-centric Methods
@@
III-C HyperEvent––A Simple Structural Encoding Baseline
@@
III-C1 Real-time Adjacency Table Maintenance ( in Figure 2)
@@
The adjacency table enables $\mathcal{O}(1)$ event retrieval, eliminating costly neighborhood searches. Moreover, the adjacency table can be rapidly obtained since its update consists of $\mathcal{O}(1)$ element addition operations, without requiring any computations, serving as a crucial foundation for the Efficient Training Algorithm introduced in Section III-D.
@@
III-C2 Context Extraction ( in Figure 2)
@@
III-C3 Relative Structural Encoding ( in Figure 2)
@@ III-C4 Classifier
III-D Efficient Training Algorithm
@@
As illustrated in Figure 4, the proposed method partitions the original event stream $\mathcal{G}$ into multiple non-overlapping segments (e.g., Event Stream 1, Event Stream 2 and Event Stream 3). Each segment corresponds to a distinct time interval demarcated by split points ($t_{1}$, $t_{a}$, $t_{b}$). Crucially, adjacency tables are precomputed at the starting timestamp of every segment, capturing the cumulative graph state up to that point. This precomputation mechanism, detailed in Section III-C1, ensures that the cumulative structural information captured at segmentation boundaries is functionally equivalent to that obtained during continuous sequential processing.
@@
TABLE I: TGB Dataset Statistics
@@
III-E Complexity Analysis
This section provides a detailed analysis of the computational efficiency of our method, examining both time and space complexity characteristics. This analysis confirms that our method achieves a highly favorable complexity profile, a necessary trait for a scalable baseline. Its efficiency stands in contrast to many more complex learning models, while its performance, as shown in Section IV, remains highly competitive.
@@ III-E Complexity Analysis
The $\mathcal{O}(|\mathcal{V}|)$ spatial scaling is substantially efficient, particularly given the multiplicative constant $n_{\text{neighbor}}$ is empirically shown to be small. When combined with the segment-level parallelism described in Section III-D, our method achieves unprecedented scalability: The framework efficiently processes temporal graphs of tens of millions of edges through parallel training, while maintaining constant-time prediction latency for individual queries, overcoming critical bottlenecks in streaming graph learning.
IV experiment
@@
IV-A Datasets
HyperEvent is evaluated using the Temporal Graph Benchmark (TGB) [47], a comprehensive repository containing five temporal graph datasets for dynamic link prediction. Following TGB’s standardized protocol, all datasets are chronologically partitioned into training (70%), validation (15%), and test sets (15%) to ensure temporally meaningful evaluation. The experiments span five large-scale datasets representing diverse real-world dynamics: 1) Wiki (tgbl-wiki-v2): Captures bipartite user-page interactions from Wikipedia co-editing logs, featuring text-attributed edges[10]. 2) Review (tgbl-review-v2): Contains weighted user-product interactions from Amazon electronics reviews (1997–2018)[48]. 3) Coin (tgbl-coin-v2): Documents cryptocurrency transactions during the 2022 Terra Luna market crash[49]. 4) Comment (tgbl-comment): Models directed reply networks from Reddit (2005–2010), comprising over 44 million edges[50]. 5) Flight (tgbl-flight-v2): Represents global airport traffic (2019–2022) with geo-attributed nodes and flight-number edges[51].
As detailed in Table I, these graphs exhibit substantial heterogeneity in scale (638K–18M nodes, 157K–67M edges), temporal dynamics (1,385–30M timestamps), and edge characteristics (weighted, directed, and attributed). The dynamic link prediction task consistently requires forecasting the next event given historical event streams.
IV-B Compared Methods
HyperEvent is compared against the following temporal graph representation learning and dynamic link prediction methods, as introduced in Section I:
@@ IV-B Compared Methods
Deep learning methods: JODIE [10], DyRep [11], TGAT [43], TGN [9], CAWN [15], NAT [14], TCL [46], GraphMixer [17], DyGFormer [16], TNCN [12], CTAN [13], TPNet [45].
IV-C Evaluation Metric
The filtered Mean Reciprocal Rank (MRR) as specified by TGB [47] is adopted, representing a ranking-based metric particularly suited for dynamic link prediction. Unlike Area Under the ROC Curve or Average Precision, which fail to adequately capture the relative ranking of positive edges against negatives, MRR directly measures the reciprocal rank of the true destination node among negative candidates. This aligns with recommendation systems and knowledge graph completion tasks where ranking quality is paramount. Crucially, all negative edges are predefined in TGB benchmark, ensuring fair comparison across methods.
@@ IV-C Evaluation Metric
IV-D Experimental Setup
TABLE II: HyperEvent Parameter Configurations on Different Datasets
@@ IV-D Experimental Setup
Dataset Configuration. All experiments implement the standardized evaluation protocols from TGB [47]. This includes chronological train/validation/test splits and standardized negative queries for dynamic link prediction. While several datasets include node features or edge attributes/weights, such information is deliberately excluded. This choice isolates and highlights HyperEvent’s capacity to capture intrinsic event interdependencies, demonstrating the framework’s efficacy without relying on external features.
HyperEvent Parameters. HyperEvent employs three key hyperparameters: $n_{\text{neighbor}}$, $n_{\text{latest}}$, and $n_{\text{segment}}$. The hyperparameter $n_{\text{segment}}$ primarily governs computational efficiency by segmenting the event stream for parallel processing. Selection of $n_{\text{segment}}$ balances GPU memory constraints with training duration relative to dataset scale. During evaluation, which involves processing multiple negative samples concurrently, a smaller $n_{\text{segment}}$ value is typically used compared to training. For the Wiki, Review, and Comment datasets, the optimal $n_{\text{neighbor}}$ was searched within {10, 15, 20}, with $n_{\text{latest}}$ fixed at 10. Observations indicated a preference for larger neighborhoods on Coin and Flight datasets, prompting a search over {10, 15, 20, 30, 40, 50} for $n_{\text{neighbor}}$. The hyperparameter $n_{\text{latest}}$ was optimized from {5, 8, 10} when fixing $n_{\text{neighbor}}=10$. Detailed hyperparameter configurations for all datasets are provided in Table II.
@@
TABLE III: MRR of methods on different datasets
@@
IV-E Methods Comparison
Table III presents the MRR of HyperEvent against 13 methods. HyperEvent demonstrates strong performance, achieving competitive results on four datasets: Wiki (80.9), Coin (77.3), Comment (75.9), and Flight (87.7), where it is outperformed only by the TPNet model. This performance is achieved through a simple framework that leverages relative structural encoding features and a lightweight transformer classifier. The fact that this straightforward approach achieves competitive results highlights the effectiveness of relative structural encoding in dynamic link prediction.

### v3

@@ Abstract
Learning representations for continuous-time dynamic graphs is critical for dynamic link prediction. While recent methods have become increasingly complex, the field lacks a strong and informative baseline to reliably gauge progress. This paper proposes HyperEvent, a simple approach that captures relative structural patterns in event sequences through an intuitive encoding mechanism. As a straightforward baseline, HyperEvent leverages relative structural encoding to identify meaningful event sequences without complex parameterization. By combining these interpretable features with a lightweight transformer classifier, HyperEvent reframes link prediction as event structure recognition. Despite its simplicity, HyperEvent achieves competitive results across multiple benchmarks, often matching the performance of more complex models. This work demonstrates that effective modeling can be achieved through simple structural encoding, providing a clear reference point for evaluating future advancements. The code for HyperEvent is accessible on GitHub(https://github.com/jianjianGJ/HyperEvent).
keywords:
continuous-time dynamic graphs , dynamic link prediction , hyper-event
journal: Knowledge-Based Systems
[…]
organization=Xidian University,city=Xi’an, postcode=710071, state=Shaanxi, country=China
{highlights}
Strong and Simple Baseline. HyperEvent provides a straightforward yet competitive baseline for dynamic link prediction, using interpretable handcrafted features and a lightweight transformer. Its simplicity helps the community distinguish genuine progress from unnecessary model complexity.
Relative Structural Encoding. The model captures meaningful event sequences through efficient relative structural encoding—measuring direct, 1-hop, and 2-hop correlations—without relying on complex learned embeddings or memory modules.
Highly Scalable and Efficient. HyperEvent supports parallelized training via segmented event streams and tensorized operations, significantly reducing training time and memory usage while maintaining competitive performance across diverse temporal graph benchmarks.
1 Introduction
@@
Temporal graphs model evolving systems through sequences of interactions (i.e., temporal events), serving as essential abstractions for real-world dynamic phenomena like social networks and transaction systems[social1, social2, trade]. These graphs are primarily modeled as discrete-time graph snapshots or continuous-time dynamic graphs (CTDGs)[tox]. This work focuses specifically on CTDGs because real-world interactions intrinsically manifest as continuous streams of events, where preserving their exact temporal sequence is critical for capturing fine-grained event relationships. Accurately predicting future events within CTDGs, known as dynamic link prediction, remains a core challenge with significant implications for applications such as real-time recommendation[recom1, recom2] and anomaly detection[anmomal1, anmomal2].
@@
Recent years have witnessed notable advances in representation learning for CTDGs. Proposed methods have aimed to address challenges such as capturing complex structural-temporal dependencies, modeling long-range interactions, and improving generalization to unseen nodes. These approaches can be broadly categorized into two paradigms: node-centric and event-centric methods. Node-centric methods maintain and update memory states for each node, using recurrent or attention-based mechanisms to evolve node embeddings over time[tgn, jodie, dyrep, tncn]. Some incorporate neural ODEs to model continuous dynamics[ctan], or employ neighbor caching to alleviate structural bottlenecks[nat]. While often scalable and inductive, these techniques reduce each event to an update of its endpoint embeddings, potentially overlooking higher-order relationships between events. In contrast, event-centric methods treat interaction events as primary entities. Techniques such as CAWN[cawn] leverage causal anonymous walks to extract temporal motifs, while DyGFormer[dygformer] uses attention mechanisms to implicitly model event group patterns. Several recent methods also highlight that relatively simple architectures—such as MLPs or patched transformers—can achieve strong performance[graphmixer, dygformer].
@@
Among these, EdgeBank[edgebank] stands out as a simple, non-parametric baseline that relies on memorizing and replaying historical interactions. However, its performance has been significantly surpassed by newer learning-based models, diminishing its utility as a meaningful benchmark. This creates a critical gap in the literature: the absence of a strong, simple baseline that can serve as a fundamental reference point for evaluating model efficacy. Without such a baseline, it becomes difficult to distinguish genuine advancements from unnecessary complexity, potentially leading to illusory progress.
@@ 1.1 Related Work
Temporal Graph Learning for Dynamic Link Prediction. Dynamic link prediction is an essential task for forecasting future interactions in continuously evolving networks, with critical applications across social networks [social, social3], communication systems [com1, com2], traffic prediction [traffic1, traffic2], and knowledge graph completion [knowledge1, knowledge2]. Unlike traditional static graph methods that rely on structural heuristics or node similarity metrics [lp1, lp2, lp3, lp4, lp5], dynamic settings require modeling temporal event streams, capturing complex dependencies among interactions, and generalizing to unseen nodes [complex1, complex2]. Early approaches discretized event streams into snapshot sequences [snap1, snap2, snap3, snap4] and applied graph neural networks [gnn1, gnn2] combined with sequence models such as RNNs [lstm, gru] or Transformers [transfomer]. However, such discretization often leads to information loss and limited expressiveness, motivating a shift toward continuous-time modeling.
@@
Continuous-Time Dynamic Graph Learning. CTDG methods can be broadly categorized into node-centric and event-centric paradigms. Node-centric approaches maintain and update memory states for each node upon every interaction. Seminal works include JODIE [jodie], which uses coupled RNNs to update user and item representations, and DyRep [dyrep], which incorporates multi-hop propagation. The Temporal Graph Network (TGN) framework unifies various memory modules and graph aggregation operators[tgn], and TGAT employs self-attention to aggregate temporal neighbor information[tgat]. Though later methods such as TNCN [tncn] and EDGE [edge] improve scalability through asynchronous updates or graph partitioning, they remain constrained by sequential, event-by-event processing that limits parallelism and overlooks higher-order event semantics. On the other hand, event-centric methods explicitly model interactions between events. CAWN [cawn] and TPNet [tpnet] encode anonymous walks to capture temporal motifs. NAT introduces reusable neighbor representations[nat], and DyGFormer applies Transformers to patched event sequences[dygformer], implicitly modeling event group patterns through attention mechanisms. TCL constructs temporal dependency graphs but relies on RNNs that struggle with long-range dependencies[tcl]. Neural ODE-based approaches like CTAN enable continuous dynamics modeling but often ignore fine-grained event correlations[ctan]. A common trend across these lines of research is the increasing architectural complexity, with many recent models incorporating sophisticated mechanisms such as attention, memory, or neural differential equations. Despite reported improvements, the field lacks a simple, yet effective baseline that clarifies the minimal complexity required for competitive performance, making it difficult to assess whether proposed enhancements genuinely capture essential temporal-topological phenomena or merely introduce inductive biases.
@@
The Role of Baselines and the Limitations of EdgeBank. Strong baselines are fundamental to meaningful scientific progress in machine learning. They serve as essential references that help the community discern actual advancements from false positives, ensure fair comparisons, and maintain a clear lower bound of performance that new methods should exceed. A poorly performing baseline can misleadingly inflate the perceived utility of new models and encourage unnecessary complexity. In temporal graph learning, EdgeBank has been adopted as a canonical baseline—a simple non-parametric method that predicts future edges based on their recurrence in historical event sequences[edgebank]. While computationally efficient and easy to implement, its performance has fallen significantly behind modern learned models across multiple datasets. This growing performance gap diminishes its utility as a meaningful reference, raising concerns about pseudo-progress in the field. When a simple heuristic is substantially outperformed, it becomes unclear whether newly proposed architectures are capturing genuine temporal dynamics or simply leveraging additional parameters and complexity to memorize dataset-specific patterns. Therefore, there is a pressing need for a new, stronger, and principled baseline that is both conceptually simple and highly competitive, enabling more rigorous evaluation and providing clearer insights into what constitutes necessary model complexity for dynamic link prediction.
2 Method
2.1 Preliminaries and Problem Formulation
@@
2.2 Existing Frameworks
2.2.1 Node-centric methods
@@
2.2.2 Event-centric Methods
@@
2.3 HyperEvent––A Simple Structural Encoding Baseline
@@
2.3.1 Real-time Adjacency Table Maintenance ( in Figure 2)
@@
The adjacency table enables $\mathcal{O}(1)$ event retrieval, eliminating costly neighborhood searches. Moreover, the adjacency table can be rapidly obtained since its update consists of $\mathcal{O}(1)$ element addition operations, without requiring any computations, serving as a crucial foundation for the Efficient Training Algorithm introduced in Section 2.4.
@@
2.3.2 Context Extraction ( in Figure 2)
@@
2.3.3 Relative Structural Encoding ( in Figure 2)
@@ 2.3.4 Classifier
2.4 Efficient Training Algorithm
@@
As illustrated in Figure 4, the proposed method partitions the original event stream $\mathcal{G}$ into multiple non-overlapping segments (e.g., Event Stream 1, Event Stream 2 and Event Stream 3). Each segment corresponds to a distinct time interval demarcated by split points ($t_{1}$, $t_{a}$, $t_{b}$). Crucially, adjacency tables are precomputed at the starting timestamp of every segment, capturing the cumulative graph state up to that point. This precomputation mechanism, detailed in Section 2.3.1, ensures that the cumulative structural information captured at segmentation boundaries is functionally equivalent to that obtained during continuous sequential processing.
@@
Table 1: TGB Dataset Statistics
@@
2.5 Complexity Analysis
This section provides a detailed analysis of the computational efficiency of our method, examining both time and space complexity characteristics. This analysis confirms that our method achieves a highly favorable complexity profile, a necessary trait for a scalable baseline. Its efficiency stands in contrast to many more complex learning models, while its performance, as shown in Section 3, remains highly competitive.
@@ 2.5 Complexity Analysis
The $\mathcal{O}(|\mathcal{V}|)$ spatial scaling is substantially efficient, particularly given the multiplicative constant $n_{\text{neighbor}}$ is empirically shown to be small. When combined with the segment-level parallelism described in Section 2.4, our method achieves unprecedented scalability: The framework efficiently processes temporal graphs of tens of millions of edges through parallel training, while maintaining constant-time prediction latency for individual queries, overcoming critical bottlenecks in streaming graph learning.
3 experiment
@@
3.1 Datasets
HyperEvent is evaluated using the Temporal Graph Benchmark (TGB) [tgb], a comprehensive repository containing five temporal graph datasets for dynamic link prediction. Following TGB’s standardized protocol, all datasets are chronologically partitioned into training (70%), validation (15%), and test sets (15%) to ensure temporally meaningful evaluation. The experiments span five large-scale datasets representing diverse real-world dynamics: 1) Wiki (tgbl-wiki-v2): Captures bipartite user-page interactions from Wikipedia co-editing logs, featuring text-attributed edges[jodie]. 2) Review (tgbl-review-v2): Contains weighted user-product interactions from Amazon electronics reviews (1997–2018)[rewiew]. 3) Coin (tgbl-coin-v2): Documents cryptocurrency transactions during the 2022 Terra Luna market crash[coin]. 4) Comment (tgbl-comment): Models directed reply networks from Reddit (2005–2010), comprising over 44 million edges[comment]. 5) Flight (tgbl-flight-v2): Represents global airport traffic (2019–2022) with geo-attributed nodes and flight-number edges[flight].
As detailed in Table 1, these graphs exhibit substantial heterogeneity in scale (638K–18M nodes, 157K–67M edges), temporal dynamics (1,385–30M timestamps), and edge characteristics (weighted, directed, and attributed). The dynamic link prediction task consistently requires forecasting the next event given historical event streams.
3.2 Compared Methods
HyperEvent is compared against the following temporal graph representation learning and dynamic link prediction methods, as introduced in Section 1:
@@ 3.2 Compared Methods
Deep learning methods: JODIE [jodie], DyRep [dyrep], TGAT [tgat], TGN [tgn], CAWN [cawn], NAT [nat], TCL [tcl], GraphMixer [graphmixer], DyGFormer [dygformer], TNCN [tncn], CTAN [ctan], TPNet [tpnet].
3.3 Evaluation Metric
The filtered Mean Reciprocal Rank (MRR) as specified by TGB [tgb] is adopted, representing a ranking-based metric particularly suited for dynamic link prediction. Unlike Area Under the ROC Curve or Average Precision, which fail to adequately capture the relative ranking of positive edges against negatives, MRR directly measures the reciprocal rank of the true destination node among negative candidates. This aligns with recommendation systems and knowledge graph completion tasks where ranking quality is paramount. Crucially, all negative edges are predefined in TGB benchmark, ensuring fair comparison across methods.
@@ 3.3 Evaluation Metric
3.4 Experimental Setup
Table 2: HyperEvent Parameter Configurations on Different Datasets
@@ 3.4 Experimental Setup
Dataset Configuration. All experiments implement the standardized evaluation protocols from TGB [tgb]. This includes chronological train/validation/test splits and standardized negative queries for dynamic link prediction. While several datasets include node features or edge attributes/weights, such information is deliberately excluded. This choice isolates and highlights HyperEvent’s capacity to capture intrinsic event interdependencies, demonstrating the framework’s efficacy without relying on external features.
HyperEvent Parameters. HyperEvent employs three key hyperparameters: $n_{\text{neighbor}}$, $n_{\text{latest}}$, and $n_{\text{segment}}$. The hyperparameter $n_{\text{segment}}$ primarily governs computational efficiency by segmenting the event stream for parallel processing. Selection of $n_{\text{segment}}$ balances GPU memory constraints with training duration relative to dataset scale. During evaluation, which involves processing multiple negative samples concurrently, a smaller $n_{\text{segment}}$ value is typically used compared to training. For the Wiki, Review, and Comment datasets, the optimal $n_{\text{neighbor}}$ was searched within {10, 15, 20}, with $n_{\text{latest}}$ fixed at 10. Observations indicated a preference for larger neighborhoods on Coin and Flight datasets, prompting a search over {10, 15, 20, 30, 40, 50} for $n_{\text{neighbor}}$. The hyperparameter $n_{\text{latest}}$ was optimized from {5, 8, 10} when fixing $n_{\text{neighbor}}=10$. Detailed hyperparameter configurations for all datasets are provided in Table 2.
@@
Table 3: MRR of methods on different datasets
@@
3.5 Methods Comparison
Table 3 presents the MRR of HyperEvent against 13 methods. HyperEvent demonstrates strong performance, achieving competitive results on four datasets: Wiki (80.9), Coin (77.3), Comment (75.9), and Flight (87.7), where it is outperformed only by the TPNet model. This performance is achieved through a simple framework that leverages relative structural encoding features and a lightweight transformer classifier. The fact that this straightforward approach achieves competitive results highlights the effectiveness of relative structural encoding in dynamic link prediction.